# Integro-differential diffusion equations on graded Lie groups

### v1

@@ Abstract.
We first study the existence, uniqueness and well-posedness of a general class of integro-differential diffusion equation on $L^{p}(\mathbb{G})$ $(1<p<+\infty$, $\mathbb{G}$ is a graded Lie group). We show the explicit solution of the considered equation by using the Fourier analysis of the group. The equation involves a nonlocal in time operator (with a general kernel) and a positive Rockland operator acting on $\mathbb{G}.$ Also, we provide $L^{p}(\mathbb{G})-L^{q}(\mathbb{G})$ $(1<p\leqslant 2\leqslant q<+\infty)$ norm estimates and time decay rate for the solutions. In fact, by using some contemporary results, one can translate the latter regularity problem to the study of boundedness of its propagator which strongly depends on the traces of the spectral projections of the Rockland operator. Moreover, in many cases, we can obtain time asymptotic decay for the solutions which depends intrinsically on the considered kernel. As a complement, we give some norm estimates for the solutions in terms of a homogeneous Sobolev space in $L^{2}(\mathbb{G})$ that involves the Rockland operator. We also give a counterpart of our results in the setting of compact Lie groups. Illustrative examples are also given.
@@
Graded Lie groups, $L^{p}$ spaces, Rockland operators, partial integro-differential equations, norm estimates, time decay rate.
@@ 1. Introduction
The Abel integral equation with kernel $k(t)=t^{\alpha-1}$ $(0<\alpha<1)$ played a very important role in the construction of the well-known family of Riemann-Liouville fractional integro-differential operators, see e.g. for some historical facts and developments [25, Chapter 1, Section 2]. Later, challenges for these type of equations came from the consideration of more general kernels. To the best of our knowledge, the first approach towards solving these equations with general kernels was by Sonine [26]. In the last years, the topic has became more popular and several researchers have been investigating some special classes of these kernels. Let us mention some of the contributions made by Clément and Nohel [4, 3], Kochubei [15], Luchko and Yamamoto [13], Vergara and Zacher [29], and others. Notice that from the point of view of evolutionary integral equations involving general classes of kernels, Prüss can serve as a reference with his book [22].
@@
whose dense (in $L^{p}(\mathbb{G})$) domain $\mathcal{D}(\mathbb{G})$ is the the space of smooth functions compactly supported in $\mathbb{G},$ see [7, Subsection 4.3.1] for more details and specifically [7, Theorem 4.3.3].
@@
Note that a kernel satisfying condition ($\mathcal{PC}$) was first considered by Sonine in [26]. Consequently, integro-differential operators with kernels $(k,l)\in\mathcal{PC}$ are usually called operators with Sonine kernels.
@@
For proving that (1.3) implies (1.1), we additionally need to assume that $l\in BV_{loc}(\mathbb{R}_{+})$ (strongly bounded variation functions).
@@
for each $v\in\mathcal{D}(\mathbb{G}).$ By [7, Theorem 4.4.3] this is the Sobolev space $L^{p}_{\nu}(\mathbb{G})$, where $\nu$ is the homogeneous order of $\mathcal{R}.$
@@
Without loss of generality, we frequently denote $u_{v}(t)$ by $u(t).$ It is known that a well-posed problem (1.3) admits a solution operator $S(t)$ (also it is called resolvent or propagator), and vice versa [22, Proposition 1.1].
@@
By [22, Theorem 4.2] and [7, Corollary 4.2.9], the integral equation (1.3) admits a resolvent in $L^{p}(\mathbb{G})$ (exponentially bounded) whenever the initial condition $u_{0}$ is continuous in $L^{p}(\mathbb{G})$. Therefore, by [22, Proposition 1.1], equation (1.3) is well-posed.
@@
Since $l\in BV_{loc}(\mathbb{R}_{+})$ we have that the resolvent is differentiable (see [22, Page 34]). So, the solution in $L^{p}(\mathbb{G})$ of problem (1.1)-(1.2) can be rewritten as $u(t)=S(t)u_{0},$ where $S(t)$ is the propagator, see e.g. [22, Chapter 2]. This form of the solution will be useful to calculate the $L^{p}(\mathbb{G})-L^{q}(\mathbb{G})$ boundedness of solutions. This will be discussed in detail in Section 3.
@@
In this work, we use the class of kernels $\mathcal{PC}$ [3, 4] which allows us to obtain a nonnegative and nonincreasing solution for the one dimensional case of our considered integro-differential equation. The latter statement will be significant and essential to prove the $L^{p}(\mathbb{G})-L^{q}(\mathbb{G})$ norm estimates for the solution of equation (1.3). Indeed, we will see that the regularity is mainly associated with the study of spectral and Fourier (Hörmander type) multipliers. In general, by using some recent results on Fourier multipliers [1, 10] that can be applied to obtain spectral multipliers, it can be showed that the norm estimates for the solution of our integro-differential difussion equation (1.1) can be reduced to the time asymptotics of its propagator in the noncommutative Lorentz space norm [16] (the space is associated with a semifinite von Neumann algebra [6, 19, 27]), which involves calculating the trace of the spectral projections of the Rockland operator $\mathcal{R}$ [23]. Moreover, we are able to establish asymptotic time behavior for the solution of equation (1.1) that will depend on the considered kernel. Notice that the first result in this direction, we mean in Lie groups, was given on compact groups in [18] and on graded Lie groups in [2], where one considered the particular case of a kernel in the form $k(t)=t^{\alpha-1}$ $(0<\alpha<1)$. Our work somehow extends previous results on $\mathbb{R}^{n}$ [29] where the boundedness of the solution of a nonlocal in time diffusion equation was provided in different spaces of $L^{p}(\mathbb{R}^{n})$ $(1<p<+\infty).$ They considered very general nonlinear-linear operators (in space) which is not possible in this paper. But, in our case, we gain generality over the space (graded Lie group) and the range of convergence for the boundedness of the solution for equation (1.1), that is extended to data in $L^{p}(\mathbb{G})$ $(1<p\leqslant 2)$ and solution in $L^{q}(\mathbb{G})$ $(2\leqslant q<+\infty).$
@@ 2.1. One dimensional integro-differential equation
It is known that (see [22, Prop. 4.5]) the assumption $(k,l)\in\mathcal{PC}$ implies that the solution of (2.1), i.e. $s_{\mu}(t)$, it is nonnegative and nonincreasing.
Next we recall some important estimates for the function $s_{\mu}(t)$ that were proved in [29, Lemma 6.1].
@@
2.2. Graded Lie groups and its Fourier analysis
Now we give the basis notions and properties of graded Lie groups that will be implicitly used through the whole paper. Full details can be found e.g. in [7, 9].
@@ 2.2. Graded Lie groups and its Fourier analysis
A family of dilations $\{D_{r}\}_{r>0}$ of a Lie algebra $\mathfrak{g}$ is a family of Lie algebra automorphisms $D_{r}:\mathfrak{g}\rightarrow\mathfrak{g}$ of the form $D_{r}=\textnormal{Exp}(A\,\textnormal{ln}r)$, where $A:\mathfrak{g}\rightarrow\mathfrak{g}$ is a diagonalisable linear operator with positive eigenvalues.
A Lie algebra that admits an automorphism of the form $D_{r}$ is nilpotent and as such it is an algebra of a connected simply connected nilpotent Lie group. Thus, any dilation $D_{r}$ gives rise to a group automorphism $D_{r}:\mathbb{G}\rightarrow\mathbb{G}$ via the global diffeomorphism $\exp_{\mathbb{G}}$ between $\mathbb{G}$ and $\mathfrak{g}$.
Let $(\pi,\mathcal{H}_{\pi})$ be a unitary representation of $\mathbb{G}$ on the separable Hilbert space $\mathcal{H}_{\pi}$. The vector $x\in\mathcal{H}_{\pi}$ is smooth if the vector valued map $x\mapsto\pi(g)x$ from $\mathbb{G}$ into $\mathcal{H}_{\pi}$ is smooth, and we denote it by $x\in\mathcal{H}^{\infty}_{\pi}$. For a strongly continuous representation $\pi$, the limit
$d\pi(X)x:=\lim_{t\rightarrow 0}\frac{1}{t}\left(\pi(\exp_{\mathbb{G}}(tX))x-x\right)\,,\quad x\in\mathcal{H}_{\pi}^{\infty},X\in\mathfrak{g}\,,$
exists in the norm topology of $\mathcal{H}_{\pi}$, and the mapping $d\pi:\mathfrak{g}\rightarrow\textnormal{End}(\mathcal{H}_{\pi}^{\infty})$ is the infinitesimal representation of $\mathfrak{g}$ on $\mathcal{H}_{\pi}^{\infty}$ associated to $\pi$. We can write $d\pi(X)=\pi(X)$.
Note that by the so-called Poincaré-Birkhoff-Witt theorem it follows that the universal enveloping algebra $\mathfrak{U}(\mathfrak{g})$ can be identified with the space of the left-invariant operators on $\mathbb{G}$. In particular we have that: if $T\in\mathfrak{U}(\mathfrak{g})$, then $T$ can be written as
$T=\sum_{\alpha\in\mathcal{J}}c_{\alpha}X^{\alpha}\,,$
where $\mathcal{J}\subset\mathbb{N}^{n}$ is a finite set, and $X^{\alpha}:=X_{1}^{\alpha_{1}}\cdots X_{n}^{\alpha_{n}}$, for $X_{j}\in\mathfrak{g}$ and $\alpha=(\alpha_{1},\cdots,\alpha_{n})\in\mathbb{N}^{n}$. This identification allows extending the domain of the mapping $d\pi:=\pi$ to $\mathfrak{U}(\mathfrak{g})$. Also, we have that for $T\in\mathfrak{U}(\mathfrak{g})$, the field of operators $\{\pi(T):\pi\in\widehat{\mathbb{G}}\}$ is the symbol in the sense of [7, Def. 5.1.33] associated to the operator $T$. By $\widehat{\mathbb{G}}$ we denote the unitary dual of $\mathbb{G}$; that is the set of all equivalence classes of irreducible, strongly continuous and unitary representations of $\mathbb{G}$.
When $\mathfrak{g}$ is a graded Lie algebra, an important class of operators in $\mathfrak{U}(\mathfrak{g})$ is the Rockland operators. These operators are frequently denoted by $\mathcal{R}$, and they are left-invariant differential operators, that are homogeneous (with respect to the dilations) of positive degree [7, Def. 3.1.15], and satisfy the Rockland condition [7, Def. 4.1.1], hence they are hypoelliptic (Let $T$ be a linear differential operator on a manifold $M$ with smooth coefficients. We say that $T$ is hypoelliptic, if for $u\in\mathcal{D}^{\prime}(N)$ the condition $Tu\in C^{\infty}(N)$ implies that $u\in C^{\infty}(N)$ for any open $N\subset M$).
Let us point out that the operator $\mathcal{R}$ on $\mathcal{D}(\mathbb{G})$ and $\pi(\mathcal{R})$ on $\mathcal{H}_{\pi}^{\infty}$ are densely defined on their domains. In the sequel, we shall keep the notation $\mathcal{R}$ and $\pi(\mathcal{R})$ for their self-adjoint extension on $L^{2}(\mathbb{G})$ and $\mathcal{H}_{\pi}^{\infty}$, respectively.
Regarding the representation of the operator $\pi(\mathcal{R})$ we have the following result: Let $\pi\in\widehat{\mathbb{G}}\setminus\{1\}$, and let $\mathcal{R}$ be a positive Rockland operator on the graded group $\mathbb{G}$. In [11] the authors proved that the spectrum of the operator $\pi(\mathcal{R})$ is discrete and lies in $(0,\infty)$. The latter allows for an orthonormal basis for $\mathcal{H}_{\pi}$ which in turn gives rise to an infinite matrix representation of the form
(2.4) $\pi(\mathcal{R})=\begin{pmatrix}\pi_{1}^{2}&0&\cdots&\cdots\\ 0&\pi_{2}^{2}&0&\cdots\\ \vdots&0&\ddots&\\ \vdots&\vdots&&\ddots\end{pmatrix}\,.$
@@
and by (2.4), the operator has the following matrix representation
@@
(2.6) $\begin{split}\int_{\mathbb{G}}|f(x)|^{2}\,{\rm d}x&=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}(\pi(f)\pi(f)^{*})\,{\rm d}\mu(\pi)\\ &=\int_{\widehat{{\mathbb{G}}}}\|\pi(f)\|^{2}_{\textnormal{HS}(\mathcal{H}_{\pi})}\,{\rm d}\mu(\pi)\,,\end{split}$
@@
(2.7) $\begin{split}f(x)&=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}(\pi(x)\pi(f))\,{\rm d}\mu(\pi)\\ &=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}(\pi(f)\pi(x))\,{\rm d}\mu(\pi)\,.\end{split}$
@@ 3.1. Well-posedness and $L^{p}-L^{q}$ estimates
Let $(k,l)\in\mathcal{PC}$ such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rockland operator of homogeneous degree $\nu$ on $\mathbb{G}$. If $u_{0}$ is a continuous function in $L^{p}(\mathbb{G})$ $(1<p<+\infty)$ then the integral equation (1.3) is well-posed. Also, the explicit solution can be written as follows:
$u(t,x)=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}[\pi(x)\overline{s}_{\pi}(t)\widehat{u_{0}}(\pi)]\,{\rm d}\mu(\pi),\quad x\in\mathbb{G}.$
@@
(3.2) $\begin{split}\partial_{t}\left(k\ast\left[\widehat{u}(t,\pi)-\widehat{u_{0}}(\pi)\right]\right)+\pi(\mathcal{R})\widehat{u}(t,\pi)&=0,\quad t>0,\\ \widehat{u}(0,\pi)&=\widehat{u_{0}}(\pi).\end{split}$
Taking into account the infinitesimal representation (2.4) of $\pi(\mathcal{R})$, the functional calculus allows the latter equation to be seen componentwise as an infinite system of equations of the form
(3.3) $\begin{split}\partial_{t}\left(k\ast\left[\widehat{u}(t,\pi)_{ij}-\widehat{u_{0}}(\pi)_{ij}\right]\right)+\pi_{i}^{2}\widehat{u}(t,\pi)_{ij}&=0,\quad t>0,\\ \widehat{u}(0,\pi)_{ij}&=\widehat{u_{0}}(\pi)_{ij},\end{split}$
where we are considering any $i,j\in\mathbb{N}$ and any $\pi\in\widehat{\mathbb{G}}.$ To solve the system given by the infinite matrix equation (3.3), we decouple the system by fixing an index $(i,j)$. Then each equation given by (3.3) consists of functions only in the time-variable $t$ and will be treated independently. Thus, normalizing equation (3.3) for fixed $(i,j)$ and taking into account that $(k,l)\in\mathcal{PC}$ we have that problem (3.3) has a nonnegative and nonincreasing solution (see subsection 2.1), which we denote by $\overline{s_{ij}}_{\pi_{i}^{2}}(t)$, and after a summation over $i,j$ with respect to the basis of $\mathcal{H}_{\pi}$, we denote it by $\overline{s}_{\pi}(t)$. Therefore, one has
@@
On the other hand, we already know that equation (1.3) is well-posed in $L^{p}(\mathbb{G})$ and the solution can be given by $u(t)=S(t)u_{0}$ (see the introduction for all details). We note that then by continuity the solution operator $u(t)=S(t)u_{0}$ to (1.3) can be extended continuously to the whole of $L^{p}(\mathbb{G})$, with values in $L^{q}(\mathbb{G}).$ Indeed, from [1, Theorem 5.1], [10, Theorem 1.1] and Lemma 2.1 it follows that
@@
Now, it was shown in [23, Theorem 8.2] that
@@
By [22, Theorem 4.2] and [7, Corollary 4.2.9], the above equation admits a resolvent in $L^{p}(\mathbb{G})$ (exponentially bounded). Thus, equation (3.5) is well-posed [22, Proposition 1.1].
@@
Let $(k,l)\in\mathcal{PC}$ such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rocklamd operator of homogeneous degree $\nu$ on $\mathbb{G}$. If $u_{0}$ is a continuous function in $L^{p}(\mathbb{G})$ $(1<p<+\infty)$, $f\in C([0,T];L^{p}(\mathbb{G}))$ and $l\ast f\in C([0,T];L^{p}(\mathbb{G}))$ then the integral equation (3.5) is well-posed. Also, if $u_{0}\in L^{p}(\mathbb{G})$ $(1<p\leqslant 2)$ and $l\ast f\in W^{1,1}([0,T];L^{p}(\mathbb{G}))$ then the following mild solution of (3.5)
@@
By [22, Proposition 1.2, item (ii)] and condition
@@
From [1, Theorem 5.1], [10, Theorem 1.1] and Lemma 2.1 it follows that

### v2

# Corresponding author

@@ Abstract.
We first study the existence, uniqueness and well-posedness of a general class of integro-differential diffusion equation on $L^{p}(\mathbb{G})$ $(1<p<+\infty$, $\mathbb{G}$ is a graded Lie group). We show the explicit solution of the considered equation. The equation involves a nonlocal in time operator (with a general kernel) and a positive Rockland operator acting on $\mathbb{G}.$ Also, we provide $L^{p}(\mathbb{G})-L^{q}(\mathbb{G})$ $(1<p\leqslant 2\leqslant q<+\infty)$ norm estimates and time decay rate for the solutions. In fact, by using some contemporary results, one can translate the latter regularity problem to the study of boundedness of its propagator which strongly depends on the traces of the spectral projections of the Rockland operator. Moreover, in many cases, we can obtain time asymptotic decay for the solutions which depends intrinsically on the considered kernel. As a complement, we give some norm estimates for the solutions in terms of a homogeneous Sobolev space in $L^{2}(\mathbb{G})$ that involves the Rockland operator. We also give a counterpart of our results in the setting of compact Lie groups. Illustrative examples are also given.
@@
Graded Lie groups, Rockland operators, partial integro-differential equations, Sonine kernel, time decay rate.
@@
The general content of the manuscript, in whole or in part, is not submitted, accepted, or published elsewhere, including conference proceedings.
@@ 1. Introduction
The Abel integral equation with kernel $k(t)=t^{\alpha-1}$ $(0<\alpha<1)$ played a very important role in the construction of the well-known family of Riemann-Liouville fractional integro-differential operators, see e.g. for some historical facts and developments [33, Chapter 1, Section 2]. Later, challenges for these type of equations came from the consideration of more general kernels. To the best of our knowledge, the first approach towards solving these equations with general kernels was by Sonine [34]. In the last years, the topic has became more popular and several researchers have been investigating some special classes of these kernels. Let us mention some of the contributions made by Clément and Nohel [5, 4], Kochubei [17], Luchko and Yamamoto [15], Vergara and Zacher [38], and others. Notice that from the point of view of evolutionary integral equations involving general classes of kernels, Prüss can serve as a reference with his book [29].
@@
whose dense (in $L^{p}(\mathbb{G})$) domain $\mathcal{D}(\mathbb{G})$ is the the space of smooth functions compactly supported in $\mathbb{G},$ see [8, Subsection 4.3.1] for more details and specifically [8, Theorem 4.3.3].
@@
It has to be highlighted that a kernel satisfying condition ($\mathcal{PC}$) was first considered by Sonine in [34]. Consequently, integro-differential operators with kernels $(k,l)\in\mathcal{PC}$ are usually called operators with Sonine kernels.
Note that the model (1.1) in $\mathbb{G}=\mathbb{R}^{n}$ with $\mathcal{R}$ being the Laplacian arises in studying the processes of anomalous and ultraslow diffusions [2, 21, 36]. From the point of view of applications in physics, researchers have focused on considering the nonlocal in time reaction-diffusion equations. In some cases, the time-fractional derivative associated to the considered problem describes the process of anomalous diffusion, dynamic processes in materials with memory, diffusion in fluids in porous media with memory, etc; we recommend to see the following sources in this direction [12, 26, 28]. More specialized works in mathematics by using this type of kernels can be found e.g. in [3, 22, 27].
Having in mind the ideas of the last paragraph and the general nature of graded Lie groups, we aim in this work to gather distinct behaviour and objects coming from the time integro-differential operator and the large class of space-operators that can be considered in this type of Lie group (e.g. the Rockland ones). In our context, an important class of problems appear when one deals with hypoelliptic operators $\mathfrak{L}$. Some of the most interesting ones are when $\mathfrak{L}$ is a Hörmander sum of squares. By Rothschild and Stein [30], one is led to considering a lifting of $\mathfrak{L}$ to a sub-Laplacian operator on a general stratified group. While, for higher order operators, one is led to considering Rockland operators.
@@
For proving that (1.3) implies (1.1), we additionally need to assume that $l\in BV_{loc}(\mathbb{R}_{+})$ (strongly bounded variation functions). It is important to highlight that everywhere below we will use some results from the book [29] where for the space $BV_{loc}(\mathbb{R}_{+})$ the author normalized the functions $k$ in this class by assuming $k(0)=0$ and $k(\cdot)$ is left-continuous on $\mathbb{R}_{+}.$ Nevertheless, in our case, it is not necessary to assume it. Also, we have verified that the additionally conditions are not necessary to use some of the results from [29].
@@
for each $v\in\mathcal{D}(\mathbb{G}).$ By [8, Theorem 4.4.3] this is the Sobolev space $L^{p}_{\nu}(\mathbb{G})$, where $\nu$ is the homogeneous order of $\mathcal{R}.$
@@
Without loss of generality, we frequently denote $u_{v}(t)$ by $u(t).$ It is known that a well-posed problem (1.3) admits a solution operator $S(t)$ (also it is called resolvent or propagator), and vice versa [29, Proposition 1.1].
@@
By [29, Theorem 4.2] and [8, Corollary 4.2.9], the integral equation (1.3) admits a resolvent in $L^{p}(\mathbb{G})$ (exponentially bounded) whenever the initial condition $u_{0}$ is continuous in $L^{p}(\mathbb{G})$. Therefore, by [29, Proposition 1.1], equation (1.3) is well-posed.
@@
Since $l\in BV_{loc}(\mathbb{R}_{+})$ we have that the resolvent is differentiable (see [29, Page 34]). So, the solution in $L^{p}(\mathbb{G})$ of problem (1.1)-(1.2) can be rewritten as $u(t)=S(t)u_{0},$ where $S(t)$ is the propagator, see e.g. [29, Chapter 2]. This form of the solution will be useful to calculate the $L^{p}(\mathbb{G})-L^{q}(\mathbb{G})$ boundedness of solutions. This will be discussed in detail in Section 3.
@@
In this work, we use the class of kernels $\mathcal{PC}$ [4, 5] which allows us to obtain a nonnegative and nonincreasing solution for the one dimensional case of our considered integro-differential equation. The latter statement will be significant and essential to prove the $L^{p}(\mathbb{G})-L^{q}(\mathbb{G})$ norm estimates for the solution of equation (1.3). Indeed, we will see that the regularity is mainly associated with the study of spectral and Fourier (Hörmander type) multipliers. In general, by using some recent results on Fourier multipliers [1, 11] that can be applied to obtain spectral multipliers, it can be showed that the norm estimates for the solution of our integro-differential difussion equation (1.1) can be reduced to the time asymptotics of its propagator in the noncommutative Lorentz space norm [18] (the space is associated with a semifinite von Neumann algebra [7, 24, 35]), which involves calculating the trace of the spectral projections of the Rockland operator $\mathcal{R}$ [31]. Moreover, we are able to establish asymptotic time behavior for the solution of equation (1.1) that will depend on the considered kernel. Notice that the first result in this direction, we mean in Lie groups, was given on compact groups in [23], where one considered the particular case of a kernel in the form $k(t)=t^{\alpha-1}$ $(0<\alpha<1)$. Our work somehow extends previous results on $\mathbb{R}^{n}$ [38] where the boundedness of the solution of a nonlocal in time diffusion equation was provided in different spaces of $L^{p}(\mathbb{R}^{n})$ $(1<p<+\infty).$
It is necessary to mention that all the results established in this paper are new in the context of graded and compact Lie groups. One the one hand, we frequently contrast our results with those already known in the classical model of $\mathbb{G}=\mathbb{R}^{n}$. In this case, usually, our results coincide with them. On the other hand, we also provide some examples of particular kernels and equations where it seems that the estimates are new in this setting and even in $\mathbb{R}^{n},$ see e.g. Section 5.
@@ 2.1. One dimensional integro-differential equation
It is known that (see [29, Prop. 4.5]) the assumption $(k,l)\in\mathcal{PC}$ implies that the solution of (2.1), i.e. $s_{\mu}(t)$, it is nonnegative and nonincreasing.
Next we recall some important estimates for the function $s_{\mu}(t)$ that were proved in [38, Lemma 6.1].
@@
2.2. Graded Lie groups and their Fourier analysis
Now we give some of the basis notions and properties of graded Lie groups that will be implicitly used through the whole paper. Full details can be found e.g. in [8, 10].
@@ 2.2. Graded Lie groups and their Fourier analysis
By $\widehat{\mathbb{G}}$ we denote the unitary dual of $\mathbb{G}$; that is the set of all equivalence classes of irreducible, strongly continuous and unitary representations of $\mathbb{G}$.
We now introduce the so-called Rockland operators. These operators are frequently denoted by $\mathcal{R}$, and they are left-invariant differential operators, that are homogeneous of positive degree [8, Def. 3.1.15], and satisfy the Rockland condition [8, Def. 4.1.1], hence they are hypoelliptic.
@@
and by using the infinitesimal representation of $\pi(\mathcal{R})$ (see [13]), the operator has the following matrix representation
@@
(2.5) $\begin{split}\int_{\mathbb{G}}|f(x)|^{2}\,{\rm d}x=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}(\pi(f)\pi(f)^{*})\,{\rm d}\mu(\pi)=\int_{\widehat{{\mathbb{G}}}}\|\pi(f)\|^{2}_{\textnormal{HS}(\mathcal{H}_{\pi})}\,{\rm d}\mu(\pi)\,,\end{split}$
@@
(2.6) $\begin{split}f(x)=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}(\pi(x)\pi(f))\,{\rm d}\mu(\pi)=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}(\pi(f)\pi(x))\,{\rm d}\mu(\pi)\,.\end{split}$
@@ 3.1. Well-posedness and $L^{p}-L^{q}$ estimates
Let $(k,l)\in\mathcal{PC}$ be such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rockland operator of homogeneous degree $\nu$ on $\mathbb{G}$. If $u_{0}$ is a continuous function in $L^{p}(\mathbb{G})$ $(1<p<+\infty)$ then the integral equation (1.3) is well-posed. Also, the explicit solution can be written as follows:
$u(t,x)=\int_{\widehat{\mathbb{G}}}\textnormal{Tr}[\pi(x)\overline{s}_{\pi}(t)\widehat{u_{0}}(\pi)]\,{\rm d}\mu(\pi),\quad x\in\mathbb{G},$
where $\overline{s}_{\pi}(t)$ is defined in the proof.
@@
(3.2) $\begin{split}\partial_{t}\left(k\ast\left[\widehat{u}(t,\pi)-\widehat{u_{0}}(\pi)\right]\right)+\pi(\mathcal{R})\widehat{u}(t,\pi)=0,\quad\widehat{u}(0,\pi)=\widehat{u_{0}}(\pi);\quad t>0.\end{split}$
Taking into account the infinitesimal representation of $\pi(\mathcal{R})$, the functional calculus allows the latter equation to be seen componentwise as an infinite system of equations of the form
(3.3) $\begin{split}\partial_{t}\left(k\ast\left[\widehat{u}(t,\pi)_{ij}-\widehat{u_{0}}(\pi)_{ij}\right]\right)+\pi_{i}^{2}\widehat{u}(t,\pi)_{ij}=0,\quad\widehat{u}(0,\pi)_{ij}=\widehat{u_{0}}(\pi)_{ij};\quad t>0.\end{split}$
where we are considering any $i,j\in\mathbb{N}$ and any $\pi\in\widehat{\mathbb{G}}.$ To solve the system given by the infinite matrix equation (3.3), we decouple the system by fixing an index $(i,j)$. Then each equation given by (3.3) consists of functions only in the time-variable $t$ and will be treated independently. Thus, normalizing equation (3.3) for fixed $(i,j)$ and taking into account that $(k,l)\in\mathcal{PC}$ we have that problem (3.3) has a nonnegative and nonincreasing solution (see Subsection 2.1), which we denote by $\overline{s_{ij}}_{\pi_{i}^{2}}(t)$, and after making it into an infinite matrix with respect to $i,j$, we denote it by $\overline{s}_{\pi}(t)$. Therefore, one has
@@
On the other hand, we already know that equation (1.3) is well-posed in $L^{p}(\mathbb{G})$ and the solution can be given by $u(t)=S(t)u_{0}$ (see the introduction for all details). We note that then by continuity the solution operator $u(t)=S(t)u_{0}$ to (1.3) can be extended continuously to the whole of $L^{p}(\mathbb{G})$, with values in $L^{q}(\mathbb{G}).$ Indeed, from [1, Theorem 5.1], [11, Theorem 1.1] and Lemma 2.1 it follows that
@@
Now, it was shown in [31, Theorem 8.2] that
@@
By [29, Theorem 4.2] and [8, Corollary 4.2.9], the above equation admits a resolvent in $L^{p}(\mathbb{G})$ (exponentially bounded). Thus, equation (3.5) is well-posed [29, Proposition 1.1].
@@
Let $(k,l)\in\mathcal{PC}$ be such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rockland operator of homogeneous degree $\nu$ on $\mathbb{G}$. If $u_{0}$ is a continuous function in $L^{p}(\mathbb{G})$ $(1<p<+\infty)$, $f\in C([0,T];L^{p}(\mathbb{G}))$ and $l\ast f\in C([0,T];L^{p}(\mathbb{G}))$ then the integral equation (3.5) is well-posed. Also, if $u_{0}\in L^{p}(\mathbb{G})$ $(1<p\leqslant 2)$ and $l\ast f\in W^{1,1}([0,T];L^{p}(\mathbb{G}))$ then the following mild solution of (3.5)
@@
By [29, Proposition 1.2, item (ii)] and condition
@@
From [1, Theorem 5.1], [11, Theorem 1.1] and Lemma 2.1 it follows that